OpenStudy (anonymous):

in an organic synthesis 62% yield of product is achieved. which of these conversions are consistent with this information 1) 74 g of butan-2-ol -> 44.64 g of butanone 2)74 g of butan-1-ol -> 54.56 g of butanoic acid 3) 74 g of 2-methylpropane-1-ol -> 54.56 g of 2-methylpropanoic acid

OpenStudy (aaronq):

You have to try all of them individually, use the equation at the bottom. Assuming that the coefficients for the reactants and products involved are 1... \(\sf moles_R=moles_P\) sub in: \(\sf moles_R=\dfrac{mass_R}{Molar~mass_R}\) \(\sf \dfrac{mass_R}{Molar~mass_R}=\dfrac{mass_P}{Molar~mass_P}\) therefore: \(\sf mass_P= \dfrac{mass_R*Molar~mass_P}{Molar~mass_R}\) this is the theoretical yield, since only 60% was made, multiply by 0.6. So, \(\sf \large mass_P= 0.6*\dfrac{mass_R*Molar~mass_P}{Molar~mass_R}\)

OpenStudy (anonymous):

For one i tried (74×72)/74 and then multiplied by 0.6 it gives 43.2. Where am i wrong?

OpenStudy (aaronq):

hm idk, i tried them too didn't get any of the possible answers

OpenStudy (anonymous):

And tomorrow's my exam :(

OpenStudy (anonymous):

Ohhhh i'm so sorry it's not 60% it's 62% my apologies

OpenStudy (aaronq):

ohh lol no worries it should be the first or the second then

OpenStudy (anonymous):

It's just i copied it from internet

OpenStudy (anonymous):

All three of them are correct according to marking scheme

OpenStudy (aaronq):

oh okay. double check them i suppose

OpenStudy (anonymous):

There must be a typo or something :/

OpenStudy (anonymous):

First two are correct the last one is creating problem now.

OpenStudy (aaronq):

i get 54.53828 for the last one

OpenStudy (anonymous):

No i got it correct yaaayyy

OpenStudy (anonymous):

It was (0.62×74×88)÷74

OpenStudy (aaronq):

awesome!